Solve the problem by using two equations with two variables:
A twin-engined aircraft can fly 864 miles from city A to city B in 3 hours with the wind and make the return trip in 8 hours against the wind. What is the speed of the wind?
I just need help setting up the equations, then i can solve fine, but word problems are not my strong point...please help.

Answer by scott8148(6628)

p is plane speed, w is wind speed

p + w = 864/3 = 288

p - w = 864/8 = 108